OpenStudy (anonymous):

I would like to get some help with: The number of state and federal inmates in millions during year x, where x > (a little line should be under this symbol) 2002 can be approximated by the following formula. N = 0.03x - 58.45 Determine the year in which there were 1.7 million inmates. Question #1: It was in the year ____________?

OpenStudy (anonymous):

Hello, anyone there??

OpenStudy (zzr0ck3r):

0.03x - 58.45=1.7 solve for x 0.03x=1.7+58.45 x = (1.7+58.45)/.03 = 2005

OpenStudy (anonymous):

x = (1.7 + 58.45) / .03 = 2005...is this a fraction ?? or does the "/" line represent division??

OpenStudy (unklerhaukus):

same thing

OpenStudy (zzr0ck3r):

yes \[a/b = \frac{a}{b}\]

OpenStudy (anonymous):

Thanks for your help!!!!

OpenStudy (zzr0ck3r):

np